help me simplify this problem
(4sqrt(3u^2/16v^8)) all one problem

Answer by dabanfield@earthlink.net(8)

(4sqrt(3u^2/16v^8)) =
4*((sqrt(3*u^2)/sqrt(16*v^(4*2))) =
4*(sqrt(3)*sqrt(u^2))/((sqrt(16)*sqrt((v^4)^2)) =
4*(sqrt(3)*u)/(4*v^4)) =
(4/4)*(sqrt(3)*u)/v^4) =
(sqrt(3)*u)/v^4